### v1

@@ 6 Proof on the gradient estimate of Stein equation
Recall the Stein equation for the distribution $\pi$, index $i$ and $1$-Lipschitz function $\phi$ is
$\Delta u(x)+\nabla\log\pi(x)\cdot\nabla u(x)=\phi(x_{i}).$
The existence of its solution is standard in elliptic theory. We will use an explicit solution by Dynkin’s formula to derive the gradient estimate. Here the main technical tool is the stochastic analysis, which can cover all the three assumptions in this paper. This extends our previous method in [8] which mainly uses the PDE method.
Let $X_{t}^{x}=X_{t}^{x}(\omega)$ be the pathwise solution of the overdamped Langevin dynamics
$\mathrm{d}X_{t}^{x}(\omega)=\nabla\log\pi(X_{t}^{x}(\omega))\mathrm{d}t+\sqrt{2}\mathrm{d}W_{t}(\omega),\quad X_{0}^{x}(\omega)=x.$ (6.1)
By Dynkin’s formula [18],
$\mathbb{E}u(X_{T}^{x})-u(x)=\mathbb{E}\int_{0}^{T}\left(\nabla\log\pi(X_{t}^{x})\cdot\nabla u(X_{t}^{x})+\Delta u(X_{t}^{x})\right)\mathrm{d}t=\int_{0}^{T}\mathbb{E}\phi(X_{t,i}^{x})\mathrm{d}t.$ (6.2)
Under any condition in 2.2, the distribution is strongly log-concave (see 2.3), and it is well-known that Law$(X_{t}^{x})$ converges to the equilibrium $\pi$ exponentially. This implies that the limit $T\rightarrow\infty$ exists for both sides in (6.2) (note $\int\phi(x_{i})\pi(x)\mathrm{d}x=0$), and the limit is
$\int u(x)\pi(x)\mathrm{d}x-u(x)=\int_{0}^{\infty}\mathbb{E}\phi(X_{t,i}^{x})\mathrm{d}t.$
Taking derivative w.r.t $x_{j}$ gives
$\nabla_{j}u(x)=-\int_{0}^{\infty}\mathbb{E}\left[\nabla_{j}X_{t,i}^{x}\cdot\nabla\phi(X_{t,i}^{x})\right]\mathrm{d}t.$
Here $\nabla_{j}X_{t,i}^{x}$ is the sample path partial derivative w.r.t $x_{j}$ and note $\omega$ is fixed. Note taking derivative on both sides is valid due to the exponential decay of $\nabla_{j}X_{t,i}^{x}$. Since $\phi$ is $1$-Lipschitz, we obtain
$\left|\nabla_{j}u(x)\right|\leq\int_{0}^{\infty}\mathbb{E}\left[\|\nabla_{j}X_{t,i}^{x}\|\left|\nabla\phi(X_{t,i}^{x})\right|\right]\mathrm{d}t\leq\mathbb{E}\int_{0}^{\infty}\|\nabla_{j}X_{t,i}^{x}\|\mathrm{d}t.$ (6.3)
So that it remains to control $\nabla_{j}X_{t,i}^{x}$. Taking derivative w.r.t. $x$ in (6.1) with $\omega$ fixed, we get
where we denote for brevity $\nabla X_{t}^{x}:=\nabla_{x}X_{t}^{x}(\omega)\in\mathbb{R}^{d\times d}$. We then apply different assumptions in 2.2 to get the gradient estimate.
$\ell_{1}$-condition
Consider the partial derivative $\nabla_{j}X_{t}^{x}\in\mathbb{R}^{d\times d_{j}}$ in (6.4)
$\mathrm{d}\nabla_{j}X_{t}^{x}=-H_{t}\cdot\nabla_{j}X_{t}^{x}\mathrm{d}t.$
We aim to control the $2$-norm of the subblock $\nabla_{j}X_{t,k}^{x}\in\mathbb{R}^{d_{k}\times d_{j}}~{}(k=1,\dots b)$. Consider fixing a test vector $v_{j}\in\mathbb{R}^{d_{j}}$ s.t. $\left|v_{j}\right|=1$, and denote $g_{t}^{x}(k,j)=\nabla_{j}X_{t,k}^{x}\cdot v_{j}\in\mathbb{R}^{d_{k}}$, then
$\frac{\mathrm{d}}{\mathrm{d}t}g_{t}^{x}(k,j)=\frac{\mathrm{d}}{\mathrm{d}t}\nabla_{j}X_{t,k}^{x}v_{j}=-H_{t}\cdot\nabla_{j}X_{t}^{x}v_{j}=-\sum_{l=1}^{b}H_{t}(k,l)\nabla_{j}X_{t,l}^{x}v_{j}=-\sum_{l=1}^{b}H_{t}(k,l)g_{t}^{x}(l,j).$
where we denote $H_{t}(k,l)\in\mathbb{R}^{d_{k}\times d_{l}}$ as the $(k,l)$-th subblock of $H_{t}$. Then by assumption,
$\frac{1}{2}\frac{\mathrm{d}}{\mathrm{d}t}\left|g_{t}^{x}(k,j)\right|^{2}=~{}$ $-\sum_{l=1}^{b}(g_{t}^{x}(k,j))^{\rm T}H_{t}(k,l)g_{t}^{x}(l,j)$
@@
Notice $\frac{1}{2}\frac{\mathrm{d}}{\mathrm{d}t}\left|g_{t}^{x}(k,j)\right|^{2}=\left|g_{t}^{x}(k,j)\right|\frac{\mathrm{d}}{\mathrm{d}t}\left|g_{t}^{x}(k,j)\right|$, we obtain
$\frac{\mathrm{d}}{\mathrm{d}t}\left|g_{t}^{x}(k,j)\right|\leq-M_{kk}\left|g_{t}^{x}(k,j)\right|+\sum_{l\neq k}M_{kl}\left|g_{t}^{x}(l,j)\right|.$
@@
$G_{0}(k,j)=|g_{0}^{x^{(j)}}(k,j)|=|\nabla_{x_{j}^{(j)}}X_{0,k}^{x^{(j)}}v_{j}|=|\nabla_{x_{j}^{(j)}}x_{k}^{(j)}v_{j}|=\delta_{jk}\left|v_{j}\right|=\delta_{jk}.$
@@
Thus we can apply A.1 and obtain
@@
$\sum_{j}|\nabla_{j}u(x^{(j)})|\leq~{}$ $\sum_{j}\int_{0}^{\infty}\mathbb{E}\|\nabla_{x^{(j)}_{j}}X_{t,i}^{x^{(j)}}\|\mathrm{d}t$
@@
Now as $x^{(j)}$ is arbitrary, we obtain the gradient estimate
$\left\|\nabla u\right\|_{\infty,1}=\sum_{j}\left\|\nabla_{j}u\right\|_{L^{\infty}}\leq c^{-1}.$
$\ell_{2}$-condition
Consider $\nabla X_{t}^{x}$ in (6.4). Denote $G_{t}=\mathrm{e}^{mt}\nabla X_{t}^{x}$ and $\tilde{H}_{t}=H_{t}-mI$, then it holds that
$\frac{\mathrm{d}}{\mathrm{d}t}G_{t}=\mathrm{e}^{mt}\left(m\nabla X_{t}^{x}-H_{t}\nabla X_{t}^{x}\right)=-\tilde{H}_{t}G_{t},\quad G_{0}=\nabla X_{0}^{x}=I_{d}.$ (6.5)
By assumption, $0\preceq\tilde{H}_{t}\preceq(M-m)I$, and $\tilde{H}_{t}$ has bandwidth $B$. By A.2, it holds that
$\mathrm{e}^{mt}\|\nabla_{j}X_{t,i}^{x}\|=\|G_{t}(j,i)\|\leq\mathrm{e}^{-(M-m)t}\sum_{k=\lceil|i-j|/B\rceil}^{\infty}\frac{t^{k}(M-m)^{k}}{k!}.$
Note this estimate holds for any initial condition $x$. For different indices $j\in[b]$, consider different initial conditions $x^{(j)}$, and take summation over $j$, we obtain
$\sum_{j}~{}$ $\|\nabla_{x^{(j)}_{j}}X_{t,i}^{x^{(j)}}\|\leq\mathrm{e}^{-mt}\mathrm{e}^{-(M-m)t}\sum_{j}\sum_{k=\lceil|i-j|/B\rceil}^{\infty}\frac{t^{k}(M-m)^{k}}{k!}$
@@
[…]
Recall (6.3), this implies
$\sum_{j}|\nabla_{j}u(x^{(j)})|\leq~{}$ $\sum_{j}\int_{0}^{\infty}\mathbb{E}\|\nabla_{x^{(j)}_{j}}X_{t,i}^{x^{(j)}}\|\mathrm{d}t=\mathbb{E}\int_{0}^{\infty}\sum_{j}\|\nabla_{x^{(j)}_{j}}X_{t,i}^{x^{(j)}}\|\mathrm{d}t$
[…]
Taking supremum over $x^{(j)}$, we obtain the gradient estimate
$\left\|\nabla u\right\|_{\infty,1}=\sum_{j}\left\|\nabla_{j}u\right\|_{L^{\infty}}\leq\frac{2BM}{m^{2}}.$
6.2 Proof of 2.7
The proof is similar to that of the $\ell_{2}$-condition case in the above section. For brevity, we only represent the different parts for the sparse graph treatment.
First note (6.3) is still valid, and we will use (6.5) to control $\nabla X_{t}^{x}$. For $G_{1}$ in (6.5), we need to modify the proof in A.2. Follow the notations therein, we can truncate the Dyson series as
Consider the off-diagonal entry $G_{1}(j,k)$. Take $n=d_{G}(j,k)-1$, then since all the path connecting $j$ and $k$ has length no less than $d_{G}(j,k)>n$, it must hold that
$\forall 1\leq q\leq n,\quad\left[(MI-H_{t_{1}})\cdots(MI-H_{t_{q}})\right](j,k)=0~{}\Rightarrow~{}\hat{X}_{q}[X](1)(j,k)=0,$
Therefore,
$G_{1}(j,k)=\exp(-M)\sum_{q=n+1}^{\infty}\frac{1}{q!}\hat{X}_{q}[X](1)(j,k).$
Note $\|\hat{X}_{q}[X](1)\|_{\rm op}\leq M^{q}$ by (A.2), we obtain
$\|G_{1}(j,k)\|\leq\exp(-M)\sum_{q=n+1}^{\infty}\frac{M^{q}}{q!}=\exp(-M)\sum_{q=d_{G}(j,k)}^{\infty}\frac{M^{q}}{q!}.$
By the same scaling argument as in the proof in A.2, we obtain
$\mathrm{e}^{mt}\|\nabla_{j}X_{t,k}^{x}\|=\|G_{t}(j,k)\|\leq\mathrm{e}^{-(M-m)t}\sum_{q=d_{G}(j,k)}^{\infty}\frac{t^{q}(M-m)^{q}}{q!}.$
$\sum_{j\in V}\|\nabla_{x^{(j)}_{j}}X_{t,k}^{x^{(j)}}\|\leq~{}$ $\mathrm{e}^{-mt}\mathrm{e}^{-(M-m)t}\sum_{j\in V}\sum_{q=d_{G}(j,k)}^{\infty}\frac{t^{q}(M-m)^{q}}{q!}$
[…]
Here we use the sparsity condition $|N_{j}^{q}\backslash\{v\}|\leq sq^{r}$. From (6.3), we obtain
$\sum_{j\in V}|\nabla_{j}u(x^{(j)})|\leq~{}$ $\sum_{j\in V}\int_{0}^{\infty}\mathbb{E}\|\nabla_{x^{(j)}_{j}}X_{t,k}^{x^{(j)}}\|\mathrm{d}t=\mathbb{E}\int_{0}^{\infty}\sum_{j\in V}\|\nabla_{x^{(j)}_{j}}X_{t,k}^{x^{(j)}}\|\mathrm{d}t$
[…]
By A.4, it holds that (denote $\kappa=\frac{M}{m}$)
$\sum_{j\in V}|\nabla_{j}u(x^{(j)})|\leq~{}$ $\frac{1}{m}+\frac{s}{M}r!\left(\frac{m}{M}\right)^{-r-1}\left(1-\frac{m}{M}\right)$
[…]
Taking supremum over $x^{(j)}$, we obtain the gradient estimate.
6.3 Proof for 2.8
The proof is based on the proof of 2.4 and 2.2. And we only represent the different part for the multiple block case.
$W_{1}(\pi_{I},\pi_{I}^{\prime})=\sup_{\phi_{0}\in{\rm Lip}_{1}(\mathbb{R}^{d_{I}})}\int\phi_{0}(x_{I})\left(\pi_{I}(x)-\pi_{I}^{\prime}(x)\right)\mathrm{d}x_{I}.$
Denote $\phi(x_{I})=\phi_{0}(x_{I})-\int\phi_{0}(x_{I})\pi^{\prime}(x_{I})\mathrm{d}x_{I}$, and let $u(x)$ solves the Stein equation
$\Delta u(x)+\nabla\log\pi^{\prime}(x)\cdot\nabla u(x)=\phi(x_{I}).$
Then using the same method as in 2.4, we obtain
$\int\phi_{0}(x_{I})\left(\pi_{I}(x)-\pi_{I}^{\prime}(x)\right)\mathrm{d}x_{I}=\sum_{j=1}^{b}\int\nabla_{j}u(x)\cdot\left(\nabla_{j}\log\pi^{\prime}(x)-\nabla_{j}\log\pi(x)\right)\pi(x)\mathrm{d}x.$
$\Rightarrow W_{1}(\pi_{I},\pi_{I}^{\prime})\leq\left\|\nabla u\right\|_{\infty,1}\cdot\max_{j}\left\|\nabla_{j}\log\pi^{\prime}-\nabla_{j}\log\pi\right\|_{L^{1}(\pi)}.$
It remains to control $\left\|\nabla u\right\|_{\infty,1}$ for the new Stein equation. The method is still the same as in Section 6.1. Note here we need to replace (6.3) by
$\left|\nabla_{j}u(x)\right|\leq\int_{0}^{\infty}\mathbb{E}\left[\|\nabla_{j}X_{t,I}^{x}\|\left|\nabla\phi(X_{t,I}^{x})\right|\right]\mathrm{d}t\leq\mathbb{E}\int_{0}^{\infty}\|\nabla_{j}X_{t,I}^{x}\|\mathrm{d}t.$
Notice $\|\nabla_{j}X_{t,I}^{x}\|\leq\sum_{i\in I}\|\nabla_{j}X_{t,i}^{x}\|$, we obtain that
$\left\|\nabla u\right\|_{\infty,1}=\sum_{j\in[b]}\left|\nabla_{j}u(x)\right|\leq\sum_{j\in[b]}\mathbb{E}\int_{0}^{\infty}\sum_{i\in I}\|\nabla_{j}X_{t,i}^{x}\|\mathrm{d}t\leq\sum_{i\in I}\delta=\delta|I|.$
This completes the proof.
@@ 7 Conclusion
In this paper, we investigate the localized structure in high-dimensional distributions, and show how it can reduce the complexity of approximating and computing the localized distributions. We introduce the $\delta$-locality condition to quantify the degree of localization, and develop a $\ell_{\infty}$-approximation theory to get a finer control of the approximation error. This theory provides new insights into the localized structure in high-dimensional distributions, and establishes a theoretical foundation for localized methods. It suggests to exploit the localized structure in algorithm design to reduce the complexity of accessing these distributions. As examples, we apply the $\ell_{\infty}$-approximation theory to the localized LIS and localized score matching method, and show how the localized structure help to reduce the computational complexity. We also show the surrogate approximation of the localized distribution does not suffer from the curse of dimensionality.
Appendix A Lemmas
Lemma A.1.
$\frac{\mathrm{d}}{\mathrm{d}t}G_{t}\leq-MG_{t},$
where $\leq$ is in the entrywise sense, and $M\in\mathbb{R}^{b\times b}$ is $\delta$-diagonal dominant with non-positive off-diagonal entries, i.e. $\forall i,j\in[b],i\neq j$,
$\sum_{j\neq i}|M_{ij}|+\delta\leq M_{ii};\quad M_{ij}\leq 0.$
Then for any $t\geq 0$, it holds
$\|G_{t}\|_{\infty}\leq\mathrm{e}^{-\delta t}\|G_{0}\|_{\infty}.$
Proof.
Denote $G_{t}^{\delta}:=\mathrm{e}^{\delta t}G_{t}$, then
$\frac{\mathrm{d}}{\mathrm{d}t}G_{t}^{\delta}=\mathrm{e}^{\delta t}\left(\frac{\mathrm{d}}{\mathrm{d}t}G_{t}+\delta G_{t}\right)\leq\mathrm{e}^{\delta t}\left(-MG_{t}+\delta G_{t}\right)=\left(-M+\delta I\right)G_{t}^{\delta}.$
$\frac{\mathrm{d}}{\mathrm{d}t}G_{t}^{\delta}\mathbf{1}\leq\left(-M+\delta I\right)G_{t}^{\delta}\mathbf{1}.$
$\forall t\geq 0,\quad|G_{t}^{\delta}\mathbf{1}|_{\infty}\leq|G^{\delta}_{0}\mathbf{1}|_{\infty},$ (A.1)
which is a reformulation of the conclusion $\left\|G_{t}\right\|_{\infty}\leq\left\|G_{0}\right\|_{\infty}$ by noticing $G_{t}^{\delta}\in\mathbb{R}_{+}^{b}$ and
$\mathrm{e}^{\delta t}\|G_{t}\|_{\infty}=\|G_{t}^{\delta}\|_{\infty}=\max_{i}\sum_{j}G_{t}^{\delta}(i,j)=|G_{t}^{\delta}\mathbf{1}|_{\infty}.$
$\frac{\mathrm{d}}{\mathrm{d}t}(G^{\delta}_{s}\mathbf{1})_{i}>0;\quad(G^{\delta}_{s}\mathbf{1})_{i}=|G^{\delta}_{s}\mathbf{1}|_{\infty}.$
On the other hand, notice by assumption on $M$,
$\frac{\mathrm{d}}{\mathrm{d}t}(G^{\delta}_{s}\mathbf{1})_{i}\leq~{}$ $\left(-M_{ii}+\delta\right)(G^{\delta}_{s}\mathbf{1})_{i}+\sum_{j\neq i}(-M_{ij})(G^{\delta}_{s}\mathbf{1})_{j}$
[…]
Contradiction. This proves our result. ∎
Lemma A.2.
Let $H_{t}\in\mathbb{R}^{d\times d}$ be a sequence of positive definite matrices with uniform block bandwidth $B>0$ and $2$-norm bound $M$, i.e.
$\forall|i-j|>B,\quad H_{t}(i,j)=0;\quad 0\preceq H_{t}\preceq MI_{d}.$
Consider the matrix ODE
$\frac{\mathrm{d}}{\mathrm{d}t}G_{t}=-H_{t}G_{t},\quad G_{0}=I_{d}.$
Then for any $t\geq 0$, it holds that
$\|G_{t}(i,j)\|\leq\exp(-tM)\sum_{k=\lceil|i-j|/B\rceil}^{\infty}\frac{(tM)^{k}}{k!}.$
Proof.
I. Scaling. First notice it suffices to consider the case $t=1$. For the general case $t=t_{0}$, consider the following scaling argument. Let $G_{t}$ be the solution, then $G_{t_{0}s}$ solves

### v2

@@ 6.1 Explicit Solution of the Stein Equation
Proof.
Similar as the proof of 2.1, it remains to control $\nabla_{x_{j}}X_{t,k}^{x}$ from (6.6). To control the $2$-norm of $\nabla_{x_{j}}X_{t,k}^{x}\in\mathbb{R}^{d_{k}\times d_{j}}~{}(k=1,\dots b)$, consider fixing a test vector $v_{j}\in\mathbb{R}^{d_{j}}$ s.t. $\left\|v_{j}\right\|=1$, and denote $g_{t}^{x}(k,j)=\nabla_{x_{j}}X_{t,k}^{x}\cdot v_{j}\in\mathbb{R}^{d_{k}}$, then
$\frac{\mathrm{d}}{\mathrm{d}t}g_{t}^{x}(k,j)=\frac{\mathrm{d}}{\mathrm{d}t}\nabla_{x_{j}}X_{t,k}^{x}v_{j}=-H_{t}\cdot\nabla_{x_{j}}X_{t}^{x}v_{j}=-\sum_{l=1}^{b}H_{t}(k,l)\nabla_{x_{j}}X_{t,l}^{x}v_{j}=-\sum_{l=1}^{b}H_{t}(k,l)g_{t}^{x}(l,j).$
Here we denote $H_{t}(k,l)\in\mathbb{R}^{d_{k}\times d_{l}}$ as the $(k,l)$-th subblock of $H_{t}$. Then by assumption,
$\frac{1}{2}\frac{\mathrm{d}}{\mathrm{d}t}\left\|g_{t}^{x}(k,j)\right\|^{2}=~{}$ $-\sum_{l=1}^{b}(g_{t}^{x}(k,j))^{\rm T}H_{t}(k,l)g_{t}^{x}(l,j)$
@@
Notice $\frac{1}{2}\frac{\mathrm{d}}{\mathrm{d}t}\left\|g_{t}^{x}(k,j)\right\|^{2}=\left\|g_{t}^{x}(k,j)\right\|\frac{\mathrm{d}}{\mathrm{d}t}\left\|g_{t}^{x}(k,j)\right\|$, we obtain
$\frac{\mathrm{d}}{\mathrm{d}t}\left\|g_{t}^{x}(k,j)\right\|\leq-M_{kk}\left\|g_{t}^{x}(k,j)\right\|+\sum_{l\neq k}M_{kl}\left\|g_{t}^{x}(l,j)\right\|.$
@@
$G_{0}(k,j)=\|g_{0}^{x^{(j)}}(k,j)\|=\|\nabla_{x_{j}^{(j)}}X_{0,k}^{x^{(j)}}v_{j}\|=\|\nabla_{x_{j}^{(j)}}x_{k}^{(j)}v_{j}\|=\delta_{jk}\left\|v_{j}\right\|=\delta_{jk}.$
@@
Thus we can apply A.3 and obtain
@@
$\sum_{j}\|\nabla_{j}u(x^{(j)})\|\leq~{}$ $\sum_{j}\int_{0}^{\infty}\mathbb{E}\|\nabla_{x^{(j)}_{j}}X_{t,i}^{x^{(j)}}\|\mathrm{d}t$
@@
Now as $x^{(j)}$ is arbitrary, we obtain the desired gradient estimate with $\delta=c^{-1}$. ∎
6.4 Proof of 3.2
Proof.
The proof is based on the proof of 2.1 and 3.1. And we only represent the different parts for the multiple block case.
$\mathsf{W}_{1}(\pi_{I},\pi_{I}^{\prime})=\sup_{\phi\in{\rm Lip}_{1}(\mathbb{R}^{d_{I}})}\int\phi(x_{I})\left(\pi_{I}(x)-\pi_{I}^{\prime}(x)\right)\mathrm{d}x_{I}.$
Given $\phi_{I}$, let $u(x)$ solves the marginal Stein equation
$\Delta u(x)+\nabla\log\pi^{\prime}(x)\cdot\nabla u(x)=\phi(x_{I})-\mathbb{E}_{\pi^{\prime}}[\phi(x_{I})].$
Similarly as in 6.1, it holds that
$\left\|\nabla_{j}u(x)\right\|\leq\int_{0}^{\infty}\mathbb{E}\left[\|\nabla_{j}X_{t,I}^{x}\|\left\|\nabla\phi(X_{t,I}^{x})\right\|\right]\mathrm{d}t\leq\int_{0}^{\infty}\mathbb{E}\|\nabla_{j}X_{t,I}^{x}\|\mathrm{d}t.$
where $X_{t}^{x}$ is the path solution of the Langevin dynamics for $\pi^{\prime}$ with initial condition $x$. Notice $\|\nabla_{j}X_{t,I}^{x}\|\leq\sum_{i\in I}\|\nabla_{j}X_{t,i}^{x}\|$, we obtain that
$\left\|\nabla u\right\|_{\infty,1}=\sum_{j\in[b]}\left\|\nabla_{j}u(x)\right\|\leq\sum_{j\in[b]}\mathbb{E}\int_{0}^{\infty}\sum_{i\in I}\|\nabla_{j}X_{t,i}^{x}\|\mathrm{d}t\leq\sum_{i\in I}\delta=\delta|I|.$
Using the same arguments in 3.1, we obtain
$\mathsf{W}_{1}(\pi_{I},\pi_{I}^{\prime})\leq~{}$ $\left\|\nabla u\right\|_{\infty,1}\cdot\max_{j}\left\|\nabla_{j}\log\pi^{\prime}-\nabla_{j}\log\pi\right\|_{L^{1}(\pi)}$
@@
This completes the proof. ∎
@@ 7 Conclusion
In this paper, we consider the locality structure in high-dimensional distributions. We introduce a quantitative analysis by proposing the $\delta$-locality condition and establishing a marginal Otto-Villani inequality. We develop the Stein’s method for analyzing the marginals, and give new characterization of the locality structure based on the Stein’s method. The marginal form of the Otto-Villani inequality provides refined approximation guarantees for low-dimensional marginals, which lays the theoretical foundation for localization methods. We also propose to apply the localization methodology in distribution approximation and sampling to design scalable algorithms in high dimensions. As examples, we show how to localize the LIS method and score matching, to design scalable algorithms for high-dimensional distributions.
It wound be interesting to extent the analysis framework to more general distributions and different locality conditions. Relaxing the log concavity assumption, and finding operations that preserve the locality would be practically interested. It is also interesting to investigate the localization methodology in other applications, such as the kernel methods and the modern generative models. We will leave these as future work.
Acknowledgements
The work of SL is partially supported by NUS Overseas Research Immersion Award (ORIA).
Appendix A Proof of Lemmas
A.1 Lemmas for 2.1
Proof of 6.2.
I. Scaling. First note by a scaling argument, it suffices to consider $t=1$. Suppose we aim to obtain the control at $t=t_{0}$. Let $G_{t}$ solves (6.5), then $G_{t_{0}s}$ solves